OpenStudy (anonymous):

The path of a diver in the water can be modeled by the funtion y= 0.16x^2-3.2x, where x is the horizontal distance in feet that the diver travels and y is the depth in feet of the diver. im not sure what the y-intercepts (roots) are but i already know what the vertex is (-16)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

i think x-intercept is the roots.....wats ur actual question ?

OpenStudy (anonymous):

oh, the question is just the first sentence with the diver in it.

OpenStudy (anonymous):

y=0.16x^2-3.2 but just dont know how to brek it dwn in numbers on my poster

OpenStudy (anonymous):

ok,now wats ur doubt.break it down as in?sorry but, i don't get wat u want to know :/

OpenStudy (anonymous):

no thanks i already got it on yahho .. wanna see?diff(.16*x^2-3.2*x, x) = 0.32*x - 3.2 solve(0.32*x - 3.2 = 0) x = 10 f(x):=( 0.16 x^2 - 3.2 x) f(10) = - 16 Vertex = (10, - 16), Minimum of f(x) When the diver travel the first 10 feet, the depth that the diver gets is 16 feet!